OpenStudy (anonymous):

find the value of x y=x-2 7y-4x=10

OpenStudy (anonymous):

7(x-2)-4x=10

OpenStudy (saifoo.khan):

x = 8 y = 6

OpenStudy (anonymous):

faster than a speeding saifoo

OpenStudy (anonymous):

lol

OpenStudy (katrinakaif):

Plug the orginal equation again 7 (x - 2) - 4x = 10 7x - 14 - 4x = 10 3x - 14 = 10 3x = 24 x = 8 y = x - 2 y = 8 - 2 y = 6

OpenStudy (saifoo.khan):

\begin{array}l\color{red}{\text{t}}\color{orange}{\text{h}}\color{#9c9a2e}{\text{a}}\color{green}{\text{t}}\color{blue}{\text{s}}\color{purple}{\text{ }}\color{purple}{\text{t}}\color{red}{\text{h}}\color{orange}{\text{e}}\color{#9c9a2e}{\text{ }}\color{green}{\text{S}}\color{blue}{\text{u}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{r}}\color{orange}{\text{ }}\color{#9c9a2e}{\text{P}}\color{green}{\text{o}}\color{blue}{\text{w}}\color{purple}{\text{e}}\color{purple}{\text{r}}\color{red}{\text{r}}\color{orange}{\text{r}}\color{#9c9a2e}{\text{r}}\color{green}{\text{r}}\color{blue}{\text{r}}\color{purple}{\text{r}}\color{purple}{\text{r}}\color{red}{\text{r}}\color{orange}{\text{r}}\color{#9c9a2e}{\text{r}}\color{green}{\text{r}}\color{blue}{\text{r}}\color{purple}{\text{r}}\color{purple}{\text{r}}\color{red}{\text{!}}\color{orange}{\text{}}\end{array}

OpenStudy (anonymous):

saifoo what calculator do u have?

OpenStudy (anonymous):

Wolfram Alpha

OpenStudy (saifoo.khan):

Casio fx-570ES

OpenStudy (katrinakaif):

Its called the brain

OpenStudy (anonymous):

computer algebra system

OpenStudy (saifoo.khan):

No, Wolfram Sucks! i hate that! >.<